# Fingering instability in dewetting capillary nanosuspensions

### v1

@@ 3.2 Capillary suspensions dewetting
where $a$ is the acceleration, $R_{bridge}$ is the bridge neck radius, $k$ is the dimensionless ratio between the initial height and diameter of the radius. Since the usual single bridge size and the maximum rupture distance between two microparticles of 10 $\mathrm{\mu}$m are 2-5 $\mathrm{\mu}$m and 2 $\mathrm{\mu}$m respectively [25], the maximum dimensionless height $k_{max}$ is 1. The resulting Reynolds number is roughly 0.01 $\ll$ 1. At small Reynolds numbers, viscous stress plays an important role during the entire stretching and the inertia effect becomes insignificant [13]. The viscous contribution of the force can be expressed as $F_{\eta}=3\pi\eta R_{bridge}^{2}u_{z}$, where the axial velocity $u_{z,max}=at$ does not exceed the substrate velocity. Since the viscosity of the bulk liquid is larger than that of the secondary liquid, the viscous contribution comes from the bulk liquid ($\approx$10-11 N) [6, 15], indicating the dominance of capillary force ($\approx$10-6N), as we reported previously [25], the illustration is as shown in Fig 6.
@@
Figure 7: Capillary suspensions without nanoparticles at different time stamps without (a) and with (b) cavitation in their final pattern, dewetted with an acceleration of 50 m/s2. The scale bars are 5 mm.
@@
The capillary suspension fingering instability (Fig. 7) resembles the suspension pattern in Fig 3. In suspension samples, inter-particle forces are limited, with microparticle movement dominated by the lifting force $F_{L}$. Once the liquid-air interfaces are created by cavitation and by the Saffman–Taylor instability from the outer encounter, numerous fine filaments are created (Fig 4c), without the attractive forces leading to emergence. For suspensions, the instability and cavitation last roughly 5 ms (Fig 3, $t=5$ to 10 ms), while the same process in capillary suspensions takes longer due to the influence of capillary forces (Fig. 7b, $t=10$ 17 ms).
When experimenting with a Newtonian fluid, the bubbles that appear in the middle can either expand quickly, leading to cavitation structures that resemble a Voronoi diagram, or they fall below the critical diameter and disappear after the stress relaxation [7, 12]. In capillary suspension case, the inter-particle capillary force can be expressed as
@@
where $\Gamma$ is the interfacial tension between the liquid phase and $\theta$ is the contact angle, which are 25.5 $\pm$ 0.7 mN/m and 50°, respectively [26, 25, 37]. The resulting toroidal bridge capillary force is about 500 nN between two spherical particles of 10 $\mathrm{\mu}$m. As fingering instability proceeds, the attractive network eventually converges in the middle and covers up the cavitation pattern (Fig. 7a, $t=15$ 17 ms), resulting in one single large bridge in the center. However, due to the loading variation caused heterogeneity inside the sample, the pressure lost can be predominately by cavitation, the created interfaces meet before the emergence into one main bridge (Fig. 7b, $t=10$ 17 ms).
Capillary suspensions show significant differences in dendritic skeleton length and branch overlap compared to pure suspensions (Supplementary Fig. S2). While cavitation typically reduces fingering length, it increases that of capillary suspensions. The total fingering length of the sample with no clear cavitation decreases by nearly 43 %, and its peak overlap decreases by 66 %. As seen from Fig 3c, cavitation meets instability and the liquid at created discontinuous interfaces retracts mostly into single dots. However, due to the strong network, the gel structures at the interfaces do not shrink, inducing a complicated pattern near the center region (Fig 7b).
3.3 Capillary nanosuspensions dewetting
In our previous study [25], we observed that incorporating even a small amount (0.01 vol%) of hydrophilic nanoparticles resulted in their uniform distribution on the surface of the microparticles. As shown in Fig 8,
@@ 3.3 Capillary nanosuspensions dewetting
these nanoparticles are proven to induce the formation of nanometer-sized liquid films. These films increase the restitution coefficient, slowing down the energy absorption upon particle collision, lubricating the relative movement, and increasing bridge instability during stretching [21, 22, 23]. As a result, the elastic component, storage modulus $G^{\prime}$ decreases by a half order of magnitude due to the elimination of Hertzian contact force, this decreases the contact force and connectivity between microparticles during stretching. The overshoot for the viscous component, the loss modulus $G^{\prime\prime}$, decreases by nearly an order of magnitude in shear rheology. Upon compression, with the bulk liquid being expelled, the average coordination number increases, and the corresponding clustering coefficient decreases, resulting in a more homogeneous network. Furthermore, nanoparticles can diminish the contact line pinning effect due to the liquid films, and the promoted liquid exchange leads to an averaged-out bridge volume [25].
Samples with nanoparticles inside the bridges lack the initial uniform shrinking stage in comparison to samples without, the fingering instability starts simultaneously with the stretching. The competition between their instability and cavitation dynamics is similar to capillary suspensions in terms of time scale. Samples of similar sizes can be cavitation-dominant or fingering-dominant without fixed criteria. However, the larger the initial diameter $D_{0}$, the more likely the cavitation develops, three samples at different time steps are as shown in supplementary Fig. S3.
@@
Figure 10: (a) Dendritic skeleton lengths are plotted against the initial sample radius for all three types of samples, the fitted power function is shown in a dashed gray line. (b) The ratio between the real dendritic skeleton length and the theoretical value. The shaded areas are to guide the eyes on data point groups. (c) Averaged overlapped branches with ROI for all three types of samples. The sample color corresponds to the legend shown in (a).
@@
The power fitted line of the fingering length can be expressed as $L_{finger}=143.6R^{1.55}$, both units are in mm. The theoretical value has a power curve fitting of smaller than 2, indicating that in general the growth in dendritic length is smaller than the growth in sample area $4\pi R^{2}$. Similar to the capillary suspension samples, the cavitation in capillary nanosuspensions also increases the fingering instability and this effect is more pronounced as the initial sample radius $R_{0}$ increases. Due to the preferred cavitation effect, the nanosuspension samples (blue) are mostly above the theoretical values. In Fig 10b, the total dendritic skeleton lengths are normalized over their initial radius, the higher the radius, the more deviated the distribution is from the fitted line. The capillary nanosuspensions are reflected by the points within the blue area, showing a positive linearity. This linearity also exists for normal capillary suspensions (red), though much less pronounced, indicated by the larger shadowed area. This narrow distribution of blue points fits perfectly with our previous observation, when nanoparticles are added to the capillary suspensions, they become structurally more homogeneous [25]. Therefore, the fingering pattern of capillary nanosuspension has a smaller sample-to-sample variation. On the other hand, the fingering ratios of the pure suspensions (gray) do not show any dependency regarding the initial sample radius, as it is more particle distribution related, the values differ from 0.7 to 1.2 randomly.
@@
In Fig. 10c, the averaged overlapped branch numbers are demonstrated over the normalized radius, despite the distinguished differences in fingering patterns, the maximal overlapped fingering numbers are all roughly 50. At the center region ($R/R_{0}<0.2$), capillary suspensions (red) have slightly lower overlapped branches than the pure suspension due to the large central bridge. The elastic network sometimes loses competition to the cavitation effect, and this is more pronounced with the presence of nanoparticles, causing the overlap number to increase. Approaching the outer part near the menisci region at the three-phase contact line ($R/R_{0}=0.6-1$), the instability of the suspension (Fig 3b and  3c) is strongly suppressed by the interparticle capillary force and the peak is shifted from $R/R_{0}=0.76$ to 0.64 With the addition of nanoparticles, the previously suppressed side branches are released approaching $R/R_{0}=0.6$, the peak of capillary nanosuspensions (blue). The discontinued branches in capillary nanosuspension, caused by lubricated particle movement, can contribute to the strong decrease in overlapped fingering numbers at the outer part.
@@
A similar problem was raised by Owens et al. [38], when Oreo cookies were split into half, adhesive failure was dominant, and nearly 95 % of the creme stayed on one wafer. This corresponds greatly to the situation of capillary suspension as shown in Fig 11a, where from the top view only 7.4 % of the sample remained. The residual is mainly focused in the sample center where cavitation occurred, from the side view, an excessive amount of the sample stayed on the bottom plate as the microparticles are internally connected, and the attractive network pulls the gel structure from the top plate during dewetting. Nanoparticles on the other hand strongly promote the detachment of microparticles from each other, resulting in a more localized deposition. Based on the eased breakup of gel strings, adhesive failure can be turned into cohesive failure with the incorporation of nanoparticles, where the samples are more evenly distributed on both plates. The summarized residual area percentages are as plotted in Fig. 11c, from which it can be concluded that the normal capillary suspensions with no nanoparticles have a very pronounced adhesion which results in poor attachment of samples on the top plate. Even for points with high residual area ratios of 20.3 % and 23.6 %, the sample failure distribution is not even. A clear preferred deposition on the top plate can be seen in Supplemental Fig. S4a and S4b respectively, with limited amounts of sample on the bottom plate. When the acceleration is tripled to 150 g/m2, this cavitation effect is much stronger, which favors the deposition as shown previously (Supplemental Fig. S5).
@@ 4 Conclusions
Microparticles in pure liquid caused increased fingers and dendritic length due to enhanced particle interaction and these inclusions serve as nucleation sites for bubbles at low pressure. By adding secondary fluid, the bridges are formed and the total fingering number decreases, with a pronounced suppression at the center and the outer region, due to the introduction of a strong sample-spanning network with high elasticity. However, due to strong network heterogeneity, their fingering instability and cavitation effect are hard to predict, the fingering length fluctuates and the final residual is uneven. With the addition of only 0.01 volnanoparticles, interesting phenomena have been observed. The capillary nanosuspensions’ overall fingering instability becomes less sample-dependent with a high linear correlation between its initial radius. This aligns with our previous study, where nanoparticles lubricate the inter-microparticle movement and homogenize the sample distribution, indicated by lower clustering coefficients in comparison to the normal capillary suspensions [25]. Meanwhile, the post-stretching residual of the gel structure can be strongly standardized by adding nanoparticles inside the liquid bridges. The elevated breakup between microparticle bridges can profoundly improve the sample deposition and change the failure type from adhesive failure, where the majority of the samples are stuck on one (random) plate, to cohesive failure, where the distribution is much more even.
@@
Lingyue Liu: Conceptualization, Methodology, Investigation, Writing – original draft. Mete Abbot: Conceptualization, Methodology, Software, Writing – original draft. Philipp Brockmann: Software, investigation, Writing – review & editing. Ilia V. Roisman: Funding acquisition, Conceptualization, Methodology, Writing – review & editing, Supervision. Jeanette Hussong: Funding acquisition, Conceptualization, Methodology, Writing – review & editing, Supervision. Erin Koos: Funding acquisition, Conceptualization, Methodology, Writing – review & editing, Supervision.

### v2

@@ 3.2 Capillary suspension dewetting
where $a$ is the acceleration, $R_{bridge}$ is the bridge neck radius, $k$ is the dimensionless ratio between the initial height and diameter of the radius. Based on our previous work [28], the single bridge size and the maximum rupture distance between two microparticles of 10 $\mathrm{\mu}$m are 2-5 $\mathrm{\mu}$m and 2 $\mathrm{\mu}$m respectively, the maximum dimensionless height $k_{max}$ is 1. The resulting Reynolds number is roughly 0.01 $\ll$ 1. At small Reynolds numbers, viscous stress plays an important role during the entire stretching and the inertia effect becomes insignificant [13]. The viscous contribution of the force can be expressed as $F_{\eta}=3\pi\eta R_{bridge}^{2}u_{z}$, where the maximum particle velocity in $z$ axis $u_{z,max}=at$, $a$ is the substrate acceleration, and $u_{z,max}$ does not exceed the substrate velocity. Since the viscosity of the bulk liquid is larger than that of the secondary liquid, the viscous contribution comes from the bulk liquid ($\approx$10-11 N) [6, 15], which is negligible in comparison to the capillary force ($\approx$10-6 N), as we reported previously [28]. As shown in Fig 6, the lift force $F_{L}$ can either be as described in Eq. 1 or $F_{cap}$, depending on whether the connecting particle is present.
@@
Figure 7: Capillary suspensions without nanoparticles at different timestamps without (a) and with (b) cavitation in their final pattern, dewetted with an acceleration of 50 m/s2. The gap height at each timestamp is $H$. The scale bars are 5 mm.
@@
The capillary suspension fingering instability (Fig. 7) resembles the suspension pattern in Fig 3. In suspension samples, interparticle forces are limited, with microparticle movement dominated by the lifting force $F_{L}$. Once the liquid-air interfaces are created by cavitation and by the Saffman–Taylor instability from the outer encounter, numerous fine filaments are created (Fig 4c), without the attractive forces leading to emergence. For suspensions, the instability and cavitation last roughly 5 ms (Fig 3, $t=5$ to 10 ms), while the same process in capillary suspensions takes longer due to the influence of capillary forces (Fig. 7b, $t=10$ 17 ms).
When experimenting with a Newtonian fluid, the bubbles that appear in the middle can either expand quickly, leading to cavitation structures that resemble a Voronoi diagram, or they fall below the critical diameter and disappear after the stress relaxation [7, 12]. In capillary suspension case, the interparticle capillary force can be expressed as
@@
where $\Gamma$ is the interfacial tension between the liquid phase and $\theta$ is the contact angle, which are 25.5 $\pm$ 0.7 mN/m and 50°, respectively [30, 28, 39]. The resulting toroidal bridge capillary force is about 500 nN between two spherical particles of 10 $\mathrm{\mu}$m. As fingering instability proceeds, the attractive network eventually converges in the middle and covers up the cavitation pattern (Fig. 7a, $t=15$ 17 ms), resulting in one single large bridge in the center, indicating an instability controlled pattern. However, samples of similar sizes can also be cavitation-dominanted due to the loading variation caused heterogeneity inside the sample. As can be seen in Fig. 7b, $t=15$ ms, the voids grow at a much faster rate at the same timestamp in Fig. 7a, and they cannot be readily compensated, resulting in a more complicated cavitation-dominated residual pattern at the center region.
Capillary suspensions show significant differences in dendritic skeleton length and intersecting branch number compared to pure suspensions (Supplementary Fig. S2). In regular suspensions, cavitation separates fingering patterns into smaller segments, which further contract due to liquid retraction. However, in capillary suspensions, the strong interparticle capillary forces maintain network stability, opposing displacement of the newly created surfaces. This internal yield stress results in an elevated fingering instability caused by cavitation. The total fingering length of the sample with no clear cavitation decreases by nearly 43 %, and its peak intersecting number decreases by 66 %. As seen from Fig. 3c, cavitation meets instability and the liquid at created discontinuous interfaces retracts mostly into single dots. However, due to the strong network, the gel structures at the interfaces do not shrink, inducing a complicated pattern near the center region and increasing the skeleton length (Fig. 7b).
3.3 Capillary nanosuspension dewetting
In our previous study where we compressed the capillary (nano)suspensions [28], we observed that incorporating even a small amount (0.01 vol%) of hydrophilic nanoparticles resulted in their uniform distribution on the surface of the microparticles. As shown in Fig 8,
@@ 3.3 Capillary nanosuspension dewetting
these nanoparticles are proven to induce the formation of nanometer-sized liquid films [40]. These films increase the restitution coefficient, slowing down the energy absorption upon particle collision, lubricating the relative movement, and increasing bridge instability during stretching [24, 25, 26]. As a result, the elastic component, storage modulus $G^{\prime}$ decreases by a half order of magnitude due to the elimination of Hertzian contact force, this decreases the contact force and connectivity between microparticles during stretching. The overshoot for the viscous component, the loss modulus $G^{\prime\prime}$, decreases by nearly an order of magnitude in shear rheology. Upon being compressed, with the bulk liquid being expelled, the average coordination number increases, and the corresponding clustering coefficient decreases, resulting in a more homogeneous network. Furthermore, nanoparticles can diminish the contact line pinning effect due to the liquid films, and the promoted liquid exchange leads to an averaged-out bridge volume [28].
Samples with nanoparticles inside the bridges lack the initial uniform shrinking stage in comparison to samples without, the fingering instability starts simultaneously with the stretching. The competition between their instability and cavitation dynamics is similar to capillary suspensions in terms of time scale. Samples of similar sizes can be cavitation-dominated or fingering-dominated without a clear pattern, reflects the inherent variability in capillary suspension microstructure. Sample preparation steps, including mixing, loading, and slight compression during experiments can influence their fingering behaviors to varying degrees, making it hard to predict precisely. Nevertheless, we observed that larger initial diameters ($D_{0}$) consistently promote cavitation development, as demonstrated by three representative samples at different timestamps in supplementary Fig. S3.
@@
Figure 10: (a) Dendritic skeleton lengths are plotted against the initial sample radius for all three types of samples, the fitted power function is shown in a dashed gray line. (b) The ratio between the real dendritic skeleton length and the theoretical value. The shaded areas are to guide the eyes on data point groups. (c) Averaged intersecting branches with ROI for all three types of samples. The sample color corresponds to the legend shown in (a).
@@
The power fitted line of the fingering length can be expressed as $L_{finger}=143.6R^{1.55}$, both units are in mm. The theoretical value has a power curve fitting of smaller than 2, indicating that in general the growth in dendritic length is smaller than the growth in sample area $4\pi R^{2}$. Similar to the capillary suspension samples, the cavitation in capillary nanosuspensions also increases the fingering instability and this effect is more pronounced as the initial sample radius $R_{0}$ increases. Due to the preferred cavitation effect, the nanosuspension samples (blue) are mostly above the theoretical values. In Fig 10b, the total dendritic skeleton lengths are normalized over their initial radius, the higher the radius, the more deviated the distribution is from the fitted line. The capillary nanosuspensions are reflected by the points within the blue area, showing a positive linearity. This linearity also exists for normal capillary suspensions (red), though much less pronounced, indicated by the larger shadowed area. This narrow distribution of blue points fits perfectly with our previous observation. When nanoparticles are added to the capillary suspensions, they enhance the relative movement of the microparticles and promote liquid exchange, thus, resulting in a more homogeneous structure, particularly when subjected to external forces [28]. Therefore, the homogeneous microstructure leads to a smaller sample-to-sample variation in the fingering pattern of capillary nanosuspension. On the other hand, the fingering ratios of the pure suspensions (gray) do not show any dependency regarding the initial sample radius, as it is more particle distribution related, the values differ from 0.7 to 1.2 randomly.
@@
In Fig. 10c, the averaged intersecting branch numbers are demonstrated over the normalized radius, despite the distinguished differences in fingering patterns, the maximal intersecting fingering numbers are all roughly 50. At the center region ($R/R_{0}<0.2$), capillary suspensions (red) have slightly lower intersecting branches than the pure suspension due to the large central bridge. The elastic network sometimes loses competition to the cavitation effect, and this is more pronounced with the presence of nanoparticles, causing the intersecting number to increase. Approaching the outer part near the menisci region at the three-phase contact line ($R/R_{0}=0.6-1$), the instability of the suspension (Fig 3b and  3c) is strongly suppressed by the interparticle capillary force and the peak is shifted from $R/R_{0}=0.76$ to 0.64 With the addition of nanoparticles, the previously suppressed side branches are released approaching $R/R_{0}=0.6$, the peak of capillary nanosuspensions (blue). The discontinued branches in capillary nanosuspension, caused by lubricated particle movement, can contribute to the strong decrease in intersecting fingering numbers at the outer part.
@@
A similar problem was raised by Owens et al. [41], when Oreo cookies were split into half, adhesive failure was dominant, and nearly 95 % of the creme stayed on one wafer. This corresponds greatly to the situation of capillary suspension as shown in Fig 11a, where from the top view only 7.4 % of the sample remained. The residual is mainly focused in the sample center where cavitation occurred, from the side view, an excessive amount of the sample stayed on the bottom plate as the microparticles are internally connected, and the attractive network pulls the gel structure from the top plate during dewetting. Nanoparticles on the other hand strongly promote the detachment of microparticles from each other, resulting in a more localized deposition. Based on the eased breakup of gel strings, adhesive failure can be turned into cohesive failure with the incorporation of nanoparticles, where the samples are more evenly distributed on both plates. The summarized residual area percentages are as plotted in Fig. 11c, from which it can be concluded that the normal capillary suspensions with no nanoparticles have a very pronounced adhesion which results in poor attachment of samples. Even for points with high residual area ratios of 20.3 % and 23.6 % on the top plate, the sample failure distribution is not even, a clear preferred deposition can be seen in Supplemental Fig. S4a and S4b respectively. When the acceleration is tripled to 150 g/m2, this cavitation effect is much stronger, which favors the deposition as shown previously (Supplemental Fig. S5).
@@ 4 Conclusions
Microparticles in pure liquid caused increased fingers and dendritic length due to enhanced particle interaction and these inclusions serve as nucleation sites for bubbles at low pressure. By adding secondary fluid, the bridges are formed and the total fingering number decreases, with a pronounced suppression at the center and the outer region, due to the introduction of a strong sample-spanning network with high elasticity. However, due to strong network heterogeneity, their fingering instability and cavitation effect are hard to predict, the fingering length fluctuates and the final residual is uneven. With the addition of only 0.01 volnanoparticles, interesting phenomena have been observed. The capillary nanosuspensions’ overall fingering instability becomes less sample-dependent with a high linear correlation between its initial radius. This aligns with our previous study, where nanoparticles lubricate the intermicroparticle movement and homogenize the sample distribution, indicated by lower clustering coefficients in comparison to the normal capillary suspensions [28]. Meanwhile, the post-stretching residual of the gel structure can be strongly standardized by adding nanoparticles inside the liquid bridges. The elevated breakup between microparticle bridges can profoundly improve the sample deposition and change the failure type from adhesive failure, where the majority of the samples are stuck on one (random) plate, to cohesive failure, where the distribution is much more even.
@@
Supporting Information
Supplementary data on the gap height increase over time, the fingering instability evolution shown through normalized area and shrinkage rate, representative dendritic skeleton length with intersecting branches, and high-speed imaging of pattern formation at different accelerations and initial sample sizes.
@@
Lingyue Liu: Conceptualization, Methodology, Investigation, Software, Writing – original draft. Mete Abbot: Conceptualization, Methodology, Investigation, Software, Writing – original draft. Philipp Brockmann: Software, investigation, Writing – review & editing. Ilia V. Roisman: Funding acquisition, Conceptualization, Methodology, Writing – review & editing, Supervision. Jeanette Hussong: Funding acquisition, Conceptualization, Methodology, Writing – review & editing, Supervision. Erin Koos: Funding acquisition, Conceptualization, Methodology, Writing – review & editing, Supervision.